pH of an aqueous solution of 25ºC is 12. On increasing its dilution, at 25ºC, upto 10 times its new pH would be :-
(1) 13 (2) 120 (3) 10 (4) 11

Grade:12th pass

1 Answers

Sandeep Bhamoo
5 years ago
Ph=14-Poh and Poh = -log(concentration of OH ion) here the concentration is decreased by a factor of 10 hence new Poh` = Poh +1 =2+1=3 hence new Ph=11 please do approve my answer if correct